the length of a rectangle is 7cm longer than width f the perimeter is 74cm find the length and the area of the rectangle
1
Expert's answer
2013-01-21T09:48:46-0500
We have that f = length - 7
and that perimeter
P = 74 =2*(f+length)
= 2*(length - 7+ length) =
= 2*(2*length -7)

2*(2*length -7) = 74
2*length - 7 =37
2*length =37+7=44
length = 44/2 =22

Therefore
width = f =length - 7 = 22-7 = 15
and thearea of the rectangle is
A = width *length = 15*22 = 330

Answer:
length = 22 cm
area = 330 cm^2